OpenStudy (anonymous):

Define the quadratic function f having x-intercepts 2 and 5, and y-intercept 5. So lost -_-

OpenStudy (whpalmer4):

You want a quadratic function that has roots at x = 2 and x = 5, because the x-intercept is simply the point at which the value of y = 0. I'll do a different set of roots, just so I'm not giving you the answer. We can write a polynomial with roots at \(x = 1\) and \(x = -3\) as \[P(x) = k(x-1)(x+3)\]because the roots are the place where the function = 0, and the product of such binomial terms will give us that behavior. The factor \(k\) is to allow us to scale the polynomial to hit the y intercept at x = 0. To find \(k\), we know that \[P(0) =k(0-1)(0+3) = 5\]\[k(-1)(3) = 5\]\[k = -\frac{5}{3}\]So for my set of roots, with a y-intercept of 5, the final polynomial would be \[P(x) = -\frac{5}{3}(x-1)(x+3) = -\frac{5}{3}(x^2+2x-3)\] Not so hard, is it?

hartnn (hartnn):

if you don't get that, ask me for an alternate way.

OpenStudy (anonymous):

thanks guys! I'll give it a shot

OpenStudy (whpalmer4):

If you don't mind, @hartnn, how about showing it to me, even if he does get it? :-)

hartnn (hartnn):

okk..

hartnn (hartnn):

let y= ax^2+bx+c i'll also take different numbers. Define the quadratic function f having x-intercepts 4 and 3, and y-intercept 2. for y-intercept , put x=0 so, y = c = 2 for x-intercept, put y=0 ax^2 +bx+2 = 0 and since u have 2 x-intercepts, u get 2 equations in a and b . a (4^2) +4b+2= 0 a (3^2) + 3b+2 =0 solve simultaneously to get a and b.

hartnn (hartnn):

i know my method is not easier, its just an alternate way.

OpenStudy (whpalmer4):

Oh, okay, sure. I did that with someone the other day to fit a parabola without even thinking about it, but of course it is equivalent. Which way would you prefer if you had say 5 or 6 roots?

hartnn (hartnn):

5 to 6 roots means 4 to 5 variables, my method will become too cumbersome/complicated to solve.

OpenStudy (whpalmer4):

Well, I'm always interested in different ways of thinking about problems, never know what is going to make the light bulb go on for any given student, thanks!